### v1

# Programming of Cellular Automata in C and C++

Abstract
This study explores running times of different ways to program cellular automata in C and C++, i.e. looping through arrays by different means, the effect of structures and objects, and the choice of data structure (array versus vector in C++) and compiler (GNU gcc versus Apple clang). The choice of data structure influenced the running time the most. The array version is more than 20-times faster than the vector version in C++. The choice of compiler also had an effect, with the GNU gcc compiler delivering 1.7-times faster programs as compared to the Apple clang compiler. Using pointers instead of array indices, using C instead of C++, and using structures and objects instead of primitive data types has little to neglectable effects on the running time. The study shows that using arrays and looping over them by index in C++ and compiled with GNU gcc reveals the best performance with respect to running time. If one is interested in multi-state cellular automata, objects can be used without loss of that performance. Future studies might investigate Apple’s Metal shader or compiler optimisation.
Index Terms:
cellular automaton, C/C++ programming, desktop computing, edge computing, large-scale simulation
I Introduction
It is known since a long time that simple programs such as elementary cellular automata can create complex patterns [1, 2]. The astonishing thing about this is that the computer program of a cellular automaton creating a simple, periodic, random, or complex pattern does not differ in complexity. Only the local update rule differs, specifically the values in a Boolean array with eight elements defining the rule. It is a surprising and important result because it shows that software does not have to be programmed in a complex way to create complex patterns or to solve complex problems. In general, and with cellular automata in particular, it should therefore be possible to create simple and small programs capable of solving complex problems in science and industry. This should even be possible to run on mobile and desktop computers, which is interesting for edge computing. Making efficient use of the ever more performant computer hardware for desktop computing, large-scale simulations can be envisioned as well.
So far, most of the software development does not make use of any of this. On the contrary, software gets more and more complex not because of the complexity of the problem to be solved but because of its ever-increasing size and number of dependencies [3] as well as the way we program it [4, 5]. Already in 1995, Niklaus Wirth pointed out in [4] that ”Software expands to fill the available memory.” and ”Software is getting slower more rapidly than hardware becomes faster.”. Some years later he also pointed out that we have lost sight of the need for carful craftsmanship in programming, even at universities [5]. Almost twenty years later, we must admit that the programming craftsmanship has almost completely died off. The industry’s appetite for feature-rich and shiny software is so big, that there is no time to develop quality software and educate programmers. Rather, high-level languages and easy-to-use frameworks with uncountable packages are used, which leads to patch-work code and to numerous dependencies [3]. The situation has become even worse with the recent advent of code generation with chatbots.
In the present paper I want to rescue some of the programming craftsmanship with respect to cellular automata although the results are applicable to other computer programs as well. Specifically, I am investigating different ways to use arrays to program cellular automata in C and C++ and compare their running time for one update iteration. Applications of cellular automata are often consisting of a large number of elements since this is a typical property of many complex systems. In addition, most applications require experimenting with different cellular automaton configurations. It is thus of great importance for applications to have practical running times.
II Methods
II-A Programs
Cellular automata can be programmed in many ways. Two of the major determinants of running time are how the cell states are represented and how they are updated according to a predefined rule. In all programs and experiments of the present study, a Boolean cell state and the following update function with rule 250 in the algebraic definition are used [2]. l, c, and r are variable names for the cell states of left, centre, and right cells, respectively.
bool update_rule250(bool l, bool c, bool r)
return l + r - l * r;
Since cellular automata divide space into cells and the state update depends on neighbouring cells, it is convenient to use arrays or a similar data structure to represent the space of a cellular automaton. The simplest way to represent a cell state is to use a primitive data type, however, one might use structs or objects to represent multiple states per cell. The questions are therefore how expensive it is to use structs or objects as well as how to loop over the cell states in the data structure to update them in each iteration. In the following, different versions in C and C++ of how to answer these questions are presented.
In all the programs, the cellular automaton size (grid size) is given as a constant integer because we usually know that and it does not change, the grid is defined as static because we want to allocate memory only once and for the lifetime of the program, and a for-loop is used to update cell states by calling the update function above.
II-A1 Program 1 in C (C Exp 1) – array of Boolean type, looped over by index
The most simple and intuitive way to program a cellular automaton in C is to use an array with elements of primitive data type Boolean and the array index to loop over it. Since the cellular automaton size is known and constant, and cells have fixed neighbouring cells in elementary cellular automata, arrays are a suitable data structure here. Since arrays are closely representing memory, it is also expected to be an efficient program.
const int grid_size = 1000001;
static bool grid_0[grid_size];
static bool grid_1[grid_size];
for(int i=1; i<grid_size-1; ++i)
grid_1[i] = update_rule250(grid_0[i-1],
grid_0[i], grid_0[i+1]);
II-A2 Program 2 in C (C Exp 2) – array of Boolean type, looped over by pointer
In C, it is also possible to access arrays with pointers instead of array subscripting [6]. Instead of using the array index to loop over the array, the second program in C uses pointers to achieve the same. After defining the Boolean grids, pointers to the start and end cells are created. In the loop, the pointers to the grids are incremented moving them from one memory location to the next.
const int grid_size = 1000001;
static bool grid_0[grid_size];
static bool grid_1[grid_size];
bool *p1_grid_0 = &grid_0[1];
bool *p2_grid_0 = &grid_0[0]+grid_size-1;
bool *p_grid_1 = &grid_1[1];
for( ; p1_grid_0 < p2_grid_0; ++p1_grid_0,
++p_grid_1)
*p_grid_1 = update_rule250(
*(p1_grid_0-1), *p1_grid_0,
*(p1_grid_0+1) );
II-A3 Program 3 in C (C Exp 3) – array of Boolean structures, looped over by index
In the present study, we are also looking at multi-state cellular automata, which means at data structures that can hold more than one state. This can be achieved with structures in C [6]. This is programmed here by first defining a structure called cell with a boolean member called state. Second, the grid array is an array of cell structures. And third, the structures in the array are accessed using the structure member operator structure.member in the for-loop.
const int grid_size = 1000001;
struct cell {
bool state;
static struct cell grid_0[grid_size];
static struct cell grid_1[grid_size];
for(int i=1; i<grid_size-1; ++i)
grid_1[i].state = update_rule250(
grid_0[i-1].state, grid_0[i].state,
grid_0[i+1].state);
II-A4 Programs 1 and 2 in C++ (C++ Exp 1 and 2) – array of Boolean type, looped over by index and pointer
Programs 1 and 2 in C++ are the same as programs 1 and 2 in C but compiled with a C++ compiler.
II-A5 Program 3 in C++ – vector of Boolean type, looped over by index
The C++ standard library provides data containers such as vector, which can hold a sequence of elements of a given type like an array, but it also allows extending the vector, it knows its size, and it can be checked against attempts to access out-of-range elements [7, 8]. These features make vector the recommended data container in C++ and is thus used in the following program. The vector with elements of Boolean type and of predefined size is defined and initialised with zero values before the for-loop in which the elements are accessed by index.
const int grid_size = 1000001;
static vector<bool> grid_0(grid_size, 0);
static vector<bool> grid_1(grid_size, 0);
for(int i=1; i<grid_size-1; ++i)
grid_1[i] = update_rule250(grid_0[i-1],
grid_0[i], grid_0[i+1]);
II-A6 Program 4 in C++ (C++ Exp 4) – vector of Boolean type, looped over by iterator
Instead of using an index, vector provides iterators that can be used to loop over it and some useful related functions [7, 8]. In program 4 in C++, iterators of type Boolean are first created. The iterator itr points to the first element of the vector and the iterator last points to the last element of the vector. For the sake of simplicity, boundaries of the cellular automata are not treated and thus itr is incremented and last is decremented. In the for-loop, the initialisation is empty because it is done before the loop already, the two iterators are compared in the condition, and the iterator is incremented in the update. Within the for-loop, push_back() is used to add the results of the update function to the new grid vector. The update function is called with values for the initial grid vector by dereferencing the iterator with *itr.
const int grid_size = 1000001;
static vector<bool> grid_0(grid_size, 0);
static vector<bool> grid_1(grid_size, 0);
vector<bool>::const_iterator itr =
grid_0.begin();
++itr;
vector<bool>::const_iterator last =
grid_0.end();
last;
for(itr; itr != last; ++itr)
grid_1.push_back(
update_rule250(*(itr-1), *itr,
*(itr+1)) );
II-A7 Program 5 in C++ (C++ Exp 5) – array of Boolean objects, looped over by index
C++ is object-oriented and thus provides classes to program multi-state cellular automata. Classes in C++ are like structures in C but in addition to member data, classes allow member functions. In program 5 in C++, a class Cell is first defined and subsequently used as a user-defined type for the grids and accessed by index and object member operator object.member in the for-loop.
const int grid_size = 1000001;
class Cell
public:
bool state;
static Cell grid_0[grid_size];
static Cell grid_1[grid_size];
for(int i=1; i<grid_size-1; ++i)
grid_1[i].state = update_rule250(
grid_0[i-1].state, grid_0[i].state,
grid_0[i+1].state);
II-B Experiments
Each of the C and C++ programs were run in experiments with corresponding numbering on a Mac Pro (2019, 2.5 GHz 28-Core Intel Xeon W, 1.5 TB 2933 MHz DDR4), each with the GNU gcc 13.2.0 and Apple clang 15.0.0 compilers. They were run for $10^{6}$, $10^{7}$, $10^{8}$, and $10^{9}$ cellular automaton grid sizes. Running time is measured using clock() of time.h because it can be used in C and C++. Only the for-loop is considered in the running time measurement.
III Results
Running times of the eight experiments for the Apple clang and GNU gcc compilers are shown in fig. 1. The top row shows the outcome for the GNU gcc compiler and the bottom row for the Apple clang compiler. The first and second columns show the results for grid size $10^{6}$ and the third column for grid size $10^{9}$. The second column omits the results for vector for better comparison of the results for arrays.
Running times scale linearly with grid size in the range from $10^{6}$ to $10^{9}$ for both compilers as shown by comparing the first and last column of fig. 1 as well as by looking at the data for grid sizes $10^{7}$ and $10^{8}$ (not shown in the present study).
The fastest programs are the once using array and index to loop over (C Experiment 1 and C++ Experiment 1), compiled with GNU gcc. The C version takes $3.222$ ms and the C++ version $3.252$ ms for the grid size $10^{6}$. The other array-based programs are, considering the variation in running time outcomes, very similar and do not make any difference from a practical point of view. While this holds for the GNU gcc compiler, the Apple clang compiler shows slightly different results and shows somewhat better results for the pointer over index with a factor $1.13$. This difference can also be observed in the middle column of fig. 1. The ratio of clang to gcc of the array and index programs is $1.71$ in C and C++.
For both compilers, vector leads to higher running times as compared to any other array program as shown in the left and right columns of fig. 1. Compiled with GNU gcc, the array (index) version is $34$ times faster than the vector (index) version and $23$ times faster than the vector (iterator) version. Compiled with Apple clang, the array (index) version is $7$ times faster than the vector (index) version and $20$ times faster than the vector (iterator) version.
Figure 1: Running times in milliseconds for different cellular automata programs in C and C++. The top row shows the results for the GNU gcc compiler and the bottom row for the Apple clang compiler. The left column shows the results for a cellular automaton grid size of $10^{6}$ and the right column of grid size $10^{9}$. The middle column shows the results without vector programs.
IV Discussion
The present study explores running times of different ways to program cellular automata in C and C++. Specifically, looping through arrays by different means, the effect of structures and objects, and the choice of data structure (array versus vector in C++) and compiler (GNU gcc versus Apple clang) were studied. The choice of data structure influenced the running time the most. The array version is more than 20-times faster than the vector version in C++. The choice of compiler also had an effect, with the GNU gcc compiler delivering 1.7-times faster programs as compared to the Apple clang compiler. Using pointers instead of array indices, using C instead of C++, and using structures and objects instead of primitive data types had little to neglectable effects on the running time.
In practice, it is highly recommended to use vector instead of array in C++ because they are safer, dynamic, more convenient, and are nevertheless efficient [7, 8]. Most developers are indeed using vector as their default data container. The present study shows that this default choice has its price in terms of running time. Due to features like dynamic size and safety, additional operations must be run in vector that a simple array does not need to run and thus this result is explainable.
It is somehow surprising that the GNU gcc provides faster programmes as compared to the Apple clang compiler on an Apple computer. One would expect Apple to optimise their compiler for their hardware. Whatever the reason might be, it is certainly the case that Apple does not focus on C or C++ and thus they might not be interested in optimising this compiler.
Regarding multi-state cellular automata or the use of structure and object, the experiments show that both are implemented very efficiently as they have hardly any effect on running time. This is an encouraging result to design programs using structures in C and classes in C++. In contrast, in C#, using objects does significantly influence running time [9].
There are limitations to this study, especially regarding its scope. It would be most interesting to investigate compiler optimisation, parallel programming, and Apple’s Metal shader. Future studies might study these aspects of programming cellular automata. In this study, I wanted to start with some fundamental aspects, which might be useful to build up in future studies.
In conclusion, using arrays and loop over them by index in C++ and compiled with GNU gcc reveals the best performance with respect to running time. If one is interested in multi-state cellular automata, objects can be used without loss of that performance. Future studies might look into Apple’s Metal shader or compiler optimisation.
Acknowledgment
The Santa Fe Institute is gratefully acknowledged for hosting the author as a visitor in Autumn 2023.